Steven, Mabel, Sofia, and Finn went on a very long trip to visit their grandfather, Max. They rode a taxi which ran a distance of 40 km within 2 hours after they left the house. During their travel, Mabel received a call from their grand uncle, Ford. His taxi had a flat wheel and asked them to pick him up. They picked him up half an hour later and resumed their way. After an hour, their taxi could not enter the woods where their grandfather lives, so, they decided all together to walk. 50 km, and 10 hours later, they reached Max’s cabin in the woods.


Plot and determine the slope of the curve.

Find the distance Steven, Mabel, Sofia, and Finn found Ford during the trip, and the distance where they started to walk.


1
Expert's answer
2020-10-07T07:27:06-0400

Assume that the velocity is constant.

They passed 40 km in 2 hours, which means that their average speed was 20 km/h.

Consider this statement: "They picked him [Ford] up half an hour later," which means that is was 0.5*20=10 km away from where Mabel received the call. By the time they took Ford for a ride, they've made 50 km total (40+10).

Now consider this: "After an hour, [...] they decided all together to walk," which means that by that time the taxi made 1*20=20 km more. So, they started walking when the distance passed by them was 50+20=70 km.